# Spectral gap bounds for reversible hybrid Gibbs chains

### v1

@@ 4.1 Random-scan Gibbs chain
This gives (14). Suppose now that $Q_{i,y_{i}}$ is positive semi-definite for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$. By (7) in Theorem 9, for $f\in L_{0}^{2}(\Pi)\setminus\{0\}$, we have
@@
The kernel $Q_{i,y_{i}}$ can be a multi-step Mtk, i.e., $Q_{i,y_{i}}(\cdot,\cdot)=\mathcal{Q}_{i,y_{i}}^{t_{i,y_{i}}}(\cdot,\cdot)$ for some positive integer $t_{i,y_{i}}$, where $\mathcal{Q}_{i,y_{i}}$ is reversible with respect to $P_{i,y_{i}}$. In this case, the bound in Corollary 10 (and in Corollaries 13 and 16 later on) holds with $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|\mathcal{Q}_{i,y_{i}}^{t_{i,y_{i}}}\|_{P_{i,y_{i}}}=\|\mathcal{Q}_{i,y_{i}}\|_{P_{i,y_{i}}}^{t_{i,y_{i}}}$, where the second equality holds because the operator $\mathcal{Q}_{i,y_{i}}$ is self-adjoint.
@@ 4.2 Data augmentation chain
Recall that $S(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,P_{2,\psi_{2}(x)}(\psi_{1}^{-1}(A))$ is the Mtk of a data augmentation chain. The chain is reversible with respect to $M_{1}$, and its convergence rate is determined by $\|S\|_{M_{1}}$. For $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, $Q_{2,y_{2}}$ is reversible with respect to $P_{2,y_{2}}$, and the corresponding chain’s convergence rate is $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$. The hybrid chain of interest is associated with the Mtk $\hat{S}(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,Q_{2,\psi_{2}(x)}(x,\psi_{1}^{-1}(A))$. By Lemma 7, $\hat{S}$ is reversible with respect to $M_{1}$. Our goal is to bound $\|\hat{S}\|_{M_{1}}$ in terms of $\|S\|_{M_{1}}$ and an upper bound on $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$. We will give two sets of bounds.
4.2.1 The first bound
We first give a comparison result involving Dirichlet forms.
@@ 4.2.1 The first bound
Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
@@
The proof of this theorem is similar to that of Theorem 9, and it will be relegated to Appendix B. Using Theorem 12, we can immediately establish the following corollary, which gives the precise form of (3).
@@
Let $f\in L_{0}^{2}(M_{1})$. Then $f\circ\psi_{1}\in L_{0}^{2}(\Pi)$, where $\circ$ denotes function composition. For $i=1,2$, let $\tilde{P}_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be defined as in (11), i.e., $\tilde{P}_{i}g(x)=\int_{\mathsf{X}}P_{i,\psi_{i}(x)}(\mathrm{d}x^{\prime})\,g(x^{\prime})$ for $x\in\mathsf{X}$. Note that $\tilde{P}_{1}(f\circ\psi_{1})=f\circ\psi_{1}$ due to the concentration of $P_{1,\psi_{1}(x)}(\cdot)$ on the set $\psi^{-1}_{1}(\psi_{1}(x))=\{x^{\prime}\in\mathsf{X}:\,\psi_{1}(x^{\prime})=\psi_{1}(x)\}$. Recall that $\tilde{P}_{1}$ and $\tilde{P}_{2}$ are orthogonal projections. Then
@@
Recall that $S$ is self-adjoint, so it is furthermore positive semi-definite. The desired result then follows from Theorem 12 through an argument that is analogous to the proof of Corollary 10. ∎
Prior to our work, the most recent result that relates $\|\hat{S}\|_{M_{1}}$ to $\|S\|_{M_{1}}$ that we know of is from Jones et al. (2014). The proof of their Theorem 6 implicitly states the following:
Proposition 14.
In the context of Example 6, suppose that there exists $C\in[0,1]$ such that, for $y_{2}\in\mathsf{Y}_{2}$, $x_{-1}\in\mathsf{Y}_{1}$, and $A\in\mathcal{F}_{1}$,
$R_{2,y_{2}}(x_{-1},A)\geq(1-C)\,\Pi_{2,y_{2}}(A).$ (17)
Then
$1-\|\hat{S}\|_{M_{1}}\geq(1-C)(1-\|S\|_{M_{1}}).$
In the context of Example 6, the inequality (17) ensures that the Markov chain associated with $Q_{2,y_{2}}$ is uniformly ergodic with convergence rate $C$ since it is equivalent to the following Doeblin’s condition:
$Q_{2,y_{2}}(x,\cdot)\geq(1-C)P_{2,y_{2}}(\cdot),\quad x\in\mathsf{X}.$
In Corollary 13, we relax this condition to $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq C$, which roughly equates to the chain associated with $Q_{2,y_{2}}$ being geometrically ergodic with convergence rate $C$ (Roberts and Rosenthal, 1997; Roberts and Tweedie, 2001).
4.2.2 The second bound
The following theorem is proved in Appendix C. It takes inspiration from Łatuszyński and Rudolf (2014), who studied the convergence rates of hybrid slice samplers, which can be regarded as hybrid data augmentation samplers targeting certain uniform distributions.
@@ 4.2.2 The second bound
for $M_{1}$-a.e. $y_{1}\in\mathsf{Y}_{1}$. If either (i) $t$ is even, or (ii) $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then, for $f\in L_{0}^{2}(M_{2})\setminus\{0\}$,
@@ 4.3 The relationship between $Q_{i,y_{i}}$ and $R_{i,y_{i}}$
Consider Examples 1 and 3, where $P_{i,y_{i}}$ is defined in terms of $\Pi_{i,y_{i}}$, and $Q_{i,y_{i}}$ is defined in terms of $R_{i,y_{i}}$. For some researchers, it may be more natural to consider the chain associated with $R_{i,y_{i}}$ than the one associated with $Q_{i,y_{i}}$. The following proposition establishes an equivalence relationship between them in terms of $L^{2}$ norm. For $i=1,\dots,n$, and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $Q_{i,y_{i}}$ can be regarded as an operator on $L_{0}^{2}(P_{i,y_{i}})$, whereas $R_{i,y_{i}}$ can be seen as an operator on $L_{0}^{2}(\Pi_{i,y_{i}})$. We draw a one-to-one correspondence between the two Hilbert spaces. A linear transformation $U:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ is an isometric isomorphism if $U$ is bijective and $\|Uf\|_{\Pi_{i,y_{i}}}=\|f\|_{P_{i,y_{i}}}$ for $f\in L_{0}^{2}(P_{i,y_{i}})$. The following result is proved in Appendix D.
@@ 5.2 A proximal sampler
Corollaries 13 and 16, along with Proposition 17, provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}=\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}$. In particular, combining the bound in Corollary 13 with Lemma 22 yields the following.
@@ 5.3 Random-scan Gibbs samplers with blocked updates
This is a special case of the random-scan Gibbs chain defined in Section 2. If we label the subsets of $[k]$ that are of cardinality $\ell$ using positive integers, then we recover the notations in Section 2. Let $\ell$ and $\ell^{\prime}$ be two block sizes such that $\ell>\ell^{\prime}$. One may then ask: is there a quantitative relationship between $\|T_{\ell}\|_{\Pi}$ and $\|T_{\ell^{\prime}}\|_{\Pi}$? This problem was considered by Qin and Wang (2023) for the case $\ell^{\prime}=\ell-1$ under the assumption that $\Pi$ admits a density function with respect to some product measure on $(\mathsf{X},\mathcal{B})$. What follows will be a direct generalization of one of their main results. We will show that the chain associated with $T_{\ell^{\prime}}$ can be viewed as a hybrid version of the Gibbs chain associated with $T_{\ell}$. Thus, one can connect $\|T_{\ell}\|_{\Pi}$ and $\|T_{\ell^{\prime}}\|_{\Pi}$ using Corollary 10 by taking $T=T_{\ell}$ and $\hat{T}=T_{\ell^{\prime}}$. Our goal is now to show that, when $\ell>\ell^{\prime}$, we can write
@@
Suppose that (A) holds. Let $\Lambda\subset[k]$ and $\Lambda^{\prime}\subset\Lambda$ be such that |Λ|=normal-$|\Lambda|=\ell$ and |Λ|=superscriptnormal-$|\Lambda^{\prime}|=\ell^{\prime}$. Then $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)=P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ for $\Pi$-a.e. $x\in\mathsf{X}$.
@@ 5.4 A hybrid slice sampler
In this scenario, the data augmentation algorithm associated with the Mtk $S(\cdot,\cdot)$ is called a slice sampler. To simulate the underlying chain, given the current state $y_{1}\in\mathsf{Y}_{1}$, one draws $y_{2}$ from the uniform distribution on $[0,\tilde{m}_{1}(y_{1})]$, then draws the next state from the uniform distribution on $G_{y_{2}}$. Regarding the spectral gap of slice samplers, the following result was recently obtained by Natarovskii et al. (2021). To state the result, for $y_{2}\in[0,\infty)$, let $v(y_{2})=\int_{\mathsf{Y}_{1}}1_{G}(y_{2},y_{1})\,\mu_{1}(\mathrm{d}y_{1})$.
@@
Corollaries 13 and 16, along with Proposition 17, provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}=\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}$. We shall use the upper bound from Corollary 13 and the lower bound from Corollary 16.
@@
By Jensen’s inequality, $\alpha_{t}\leq\beta_{t}$, so the lower bound in (20) is slightly sharper than that in (21). The reader can refer to Łatuszyński and Rudolf (2014) for some choices of practically useful $R_{2,y_{2}}$ with expressions for $\beta_{t}$.

### v2

@@ 4.1 Random-scan Gibbs chain
This gives (15).
Suppose now that $Q_{i,y_{i}}$ is positive semi-definite for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$. By (8) in Theorem 10, for $f\in L_{0}^{2}(\Pi)\setminus\{0\}$, we have
@@
The kernel $Q_{i,y_{i}}$ can be a multi-step Mtk, i.e., $Q_{i,y_{i}}(\cdot,\cdot)=\mathcal{Q}_{i,y_{i}}^{t_{i,y_{i}}}(\cdot,\cdot)$ for some positive integer $t_{i,y_{i}}$, where $\mathcal{Q}_{i,y_{i}}$ is reversible with respect to $P_{i,y_{i}}$. In this case, the bound in Corollary 11 (and in Corollaries 14 and 16 later on) holds with $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|\mathcal{Q}_{i,y_{i}}^{t_{i,y_{i}}}\|_{P_{i,y_{i}}}=\|\mathcal{Q}_{i,y_{i}}\|_{P_{i,y_{i}}}^{t_{i,y_{i}}}$, where the second equality holds because the operator $\mathcal{Q}_{i,y_{i}}$ is self-adjoint.
@@ 4.2 Data augmentation chain
Recall that $S(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,P_{2,\psi_{2}(x)}(\psi_{1}^{-1}(A))$ is the Mtk of a data augmentation chain. The chain is reversible with respect to $M_{1}$, and its convergence rate is determined by $\|S\|_{M_{1}}$. For $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, $Q_{2,y_{2}}$ is reversible with respect to $P_{2,y_{2}}$, and the corresponding chain’s convergence rate is $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$. The hybrid chain of interest is associated with the Mtk $\hat{S}(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,Q_{2,\psi_{2}(x)}(x,\psi_{1}^{-1}(A))$. By Lemma 7, $\hat{S}$ is reversible with respect to $M_{1}$. Our goal is to bound $\|\hat{S}\|_{M_{1}}$ in terms of $\|S\|_{M_{1}}$ and an upper bound on $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$.
4.2.1 Andrieu et al.’s (2018) bound
We first state a bound from Andrieu et al. (2018). For $y_{2}\in\mathsf{Y}_{2}$, let $L_{0}^{2}(P_{2,y_{2};\psi_{1}})$ be the set of functions $g\in L_{0}^{2}(P_{2,y_{2}})$ such that $g(x)$ depends on $x$ only through $\psi_{1}(x)$. Assume for now that
$y_{2}\mapsto\vartheta(y_{2}):=\sup_{L_{0}^{2}(P_{2,y_{2};\psi_{1}})\setminus\{0\}}\frac{\langle f,Q_{2,y_{2}}f\rangle_{P_{2,y_{2}}}}{\|f\|_{P_{2,y_{2}}}^{2}}$
is measurable. Let
$\varrho=1-\sup_{f\in L_{0}^{2}(M_{1})}\frac{\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\vartheta(y_{2})\,\|f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})\|_{P_{2,y_{2}}}^{2}}{\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\|f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})\|_{P_{2,y_{2}}}^{2}},$
where $\circ$ means function composition. Andrieu et al. (2018) showed that, for $f\in L_{0}^{2}(M_{1})$,
$\varrho(1-\langle f,Sf\rangle_{M_{1}})\leq 1-\langle f,\hat{S}f\rangle_{M_{1}}\leq 2(1-\langle f,Sf\rangle_{M_{1}}).$
The quantity $\varrho$ is apparently intractable in the sense that there is no clear way to compute it even if $y_{2}\mapsto\vartheta(y_{2})$ is known. But as noted by Andrieu et al. (2018) themselves, one can at least bound $\varrho$ from below by $1-\sup_{y_{2}}\vartheta(y_{2})$. Using this fact and making some adjustments to Andrieu et al.’s (2018) argument, we can arrive at the following theorem.
@@
(Andrieu et al., 2018, Theorem 25, Remark 26) Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
@@
For completeness, we will give the proof of Theorem 13 in Appendix B.
Using Theorem 13, we can immediately establish the following corollary, which gives the precise form of (4).
@@
Let $f\in L_{0}^{2}(M_{1})$. Then $f\circ\psi_{1}\in L_{0}^{2}(\Pi)$. For $i=1,2$, let $\tilde{P}_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be defined as in (12), i.e., $\tilde{P}_{i}g(x)=\int_{\mathsf{X}}P_{i,\psi_{i}(x)}(\mathrm{d}x^{\prime})\,g(x^{\prime})$ for $x\in\mathsf{X}$. Note that $\tilde{P}_{1}(f\circ\psi_{1})=f\circ\psi_{1}$ due to the concentration of $P_{1,\psi_{1}(x)}(\cdot)$ on the set $\psi^{-1}_{1}(\psi_{1}(x))=\{x^{\prime}\in\mathsf{X}:\,\psi_{1}(x^{\prime})=\psi_{1}(x)\}$. Recall that $\tilde{P}_{1}$ and $\tilde{P}_{2}$ are orthogonal projections. Then
@@
Recall that $S$ is self-adjoint, so it is furthermore positive semi-definite.
The desired result then follows from Theorem 13 through an argument that is analogous to the proof of Corollary 11. ∎
Note the similarity between Theorems 10 and 13, and between Corollaries 11 and 14. One undesirable feature of these results is that the bounds they produce are useful only we have nontrivial uniform bounds on quantities like $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}$. For instance, in Corollary 14, the lower bound on $1-\|\hat{S}\|_{M_{1}}$ is nontrivial (i.e., strictly greater than 0) only if $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$ is uniformly bounded away from unity. The main goal of Section 4.2 is to relax this requirement. It turns out that, for hybrid data augmentation algorithms, we can achieve this through an argument from Łatuszyński and Rudolf (2014), who studied hybrid slice samplers. (We were not able to make a meaningful improvement on Theorem 10 and Corollary 11 through an analogous treatment.)
4.2.2 Relaxing the uniform boundedness condition
The following theorem is proved in Appendix C. Loosely speaking, the proof involves analyzing the intractable quantity $\varrho$ in Andrieu et al.’s (2018) original bound using techniques from Łatuszyński and Rudolf (2014).
@@ 4.2.2 Relaxing the uniform boundedness condition
for $M_{1}$-a.e. $y_{1}\in\mathsf{Y}_{1}$. If either (i) $t$ is even, or (ii) $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then, for $f\in L_{0}^{2}(M_{1})\setminus\{0\}$,
@@
As mentioned in Section 2.2, when $\|S\|_{M_{1}}<1$, for the lower bound on $1-\|\hat{S}\|_{M_{1}}^{t}$ in Corollary 16 to be nontrivial for large values of $t$, it suffices to have $\alpha_{t}\to 0$ as $t\to\infty$. Instead of requiring that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$ is bounded away from 1 on $\mathsf{Y}_{2}$ like the bound in Corollary 14 does, here the requirement is only that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$ is bounded away from 1 on some subset of $\mathsf{Y}_{2}$, and the mass of that subset evaluated by $P_{1,y_{1}}\circ\psi_{2}^{-1}(\cdot)$ is bounded away from 0 as $y_{1}$ varies in $\mathsf{Y}_{1}$.
@@ 4.3 The relationship between $Q_{i,y_{i}}$ and $R_{i,y_{i}}$
Consider Examples 1 and 3, where $P_{i,y_{i}}$ is defined in terms of $\Pi_{i,y_{i}}$, and $Q_{i,y_{i}}$ is defined in terms of $R_{i,y_{i}}$. For some researchers, it may be more natural to consider the chain associated with $R_{i,y_{i}}$ than the one associated with $Q_{i,y_{i}}$. The following proposition establishes an equivalence relationship between them in terms of $L^{2}$ norm.
For $i=1,\dots,n$, and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $Q_{i,y_{i}}$ can be regarded as an operator on $L_{0}^{2}(P_{i,y_{i}})$, whereas $R_{i,y_{i}}$ can be seen as an operator on $L_{0}^{2}(\Pi_{i,y_{i}})$. We draw a one-to-one correspondence between the two Hilbert spaces. A linear transformation $U:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ is an isometric isomorphism if $U$ is bijective and $\|Uf\|_{\Pi_{i,y_{i}}}=\|f\|_{P_{i,y_{i}}}$ for $f\in L_{0}^{2}(P_{i,y_{i}})$.
The following result is proved in Appendix D.
@@ 5.2 A proximal sampler
Corollaries 14 and 16, along with Proposition 17, provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}=\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}$. In particular, We may use Andrieu et al.’s (2018) bound from Corollary 14 in conjunction with Lemma 22 to derive the following.
@@ 5.3 Random-scan Gibbs samplers with blocked updates
This is a special case of the random-scan Gibbs chain defined in Section 2. If we label the subsets of $[k]$ that are of cardinality $\ell$ using positive integers, then we recover the notations in Section 2.
Let $\ell$ and $\ell^{\prime}$ be two block sizes such that $\ell>\ell^{\prime}$. One may then ask: is there a quantitative relationship between $\|T_{\ell}\|_{\Pi}$ and $\|T_{\ell^{\prime}}\|_{\Pi}$? This problem was considered by Qin and Wang (2023) for the case $\ell^{\prime}=\ell-1$ under the assumption that $\Pi$ admits a density function with respect to some product measure on $(\mathsf{X},\mathcal{B})$. What follows will be a direct generalization of one of their main results. We will show that the chain associated with $T_{\ell^{\prime}}$ can be viewed as a hybrid version of the Gibbs chain associated with $T_{\ell}$. Thus, one can connect $\|T_{\ell}\|_{\Pi}$ and $\|T_{\ell^{\prime}}\|_{\Pi}$ using Corollary 11 by taking $T=T_{\ell}$ and $\hat{T}=T_{\ell^{\prime}}$.
Our goal is now to show that, when $\ell>\ell^{\prime}$, we can write
@@
Suppose that (A) holds. Let $\Lambda\subset[k]$ and $\Lambda^{\prime}\subset\Lambda$ be such that $|\Lambda|=\ell$ and $|\Lambda^{\prime}|=\ell^{\prime}$. Then $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)=P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ for $\Pi$-a.e. $x\in\mathsf{X}$.
@@ 5.4 A hybrid slice sampler
In this scenario, the data augmentation algorithm associated with the Mtk $S(\cdot,\cdot)$ is called a slice sampler. To simulate the underlying chain, given the current state $y_{1}\in\mathsf{Y}_{1}$, one draws $y_{2}$ from the uniform distribution on $[0,\tilde{m}_{1}(y_{1})]$, then draws the next state from the uniform distribution on $G_{y_{2}}$.
Regarding the spectral gap of slice samplers, the following result was recently obtained by Natarovskii et al. (2021). To state the result, for $y_{2}\in[0,\infty)$, let $v(y_{2})=\int_{\mathsf{Y}_{1}}1_{G}(y_{2},y_{1})\,\mu_{1}(\mathrm{d}y_{1})$.
@@
Corollaries 14 and 16, along with Proposition 17, provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}=\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}$. We shall use the upper bound from Corollary 14 and the lower bound from Corollary 16.
@@
By Jensen’s inequality, $\alpha_{t}\leq\beta_{t}$, so the lower bound in (20) is slightly sharper than that in (21).
The reader can refer to Łatuszyński and Rudolf (2014) for some choices of practically useful $R_{2,y_{2}}$ with expressions for $\beta_{t}$.
After the initial release of our work, Power et al. (2024) established some new convergence bounds for exact and hybrid slice samplers that further extended Proposition 28.